OpenStudy (anonymous):

Given that x^(2)+x=1, find the value of x^(3)+2x^(2)+2

myininaya (myininaya):

Have you solve \[x^2+x-1=0\] yet?

OpenStudy (bahrom7893):

Solve for x

OpenStudy (anonymous):

I'm supposed to modify one of the equations such that the other one can be substituted in

myininaya (myininaya):

Solve the equation I mentioned and we will get to evaluating the expression for two values of x that we get

OpenStudy (anonymous):

X=0.6180339887 OR X=-1.618033989

myininaya (myininaya):

Can you approximate?

myininaya (myininaya):

I mean are you allowed?

OpenStudy (mertsj):

Well if that is the case, you could say that x = 1-x^2 and substitute that into the second equation but you will not get a numerical answer.

myininaya (myininaya):

\[x=\frac{-1 \pm \sqrt{1-4(1)(-1)}}{2}=\frac{-1 \pm \sqrt{5}}{2}= \frac{-1 + \sqrt{5}}{2} \text{ or } \frac{-1 - \sqrt{5}}{2}\]

OpenStudy (anonymous):

I'm not sure... I tried to substitute these two values into the second equation in the question but I got 3 & 3

OpenStudy (anonymous):

How do I modify the 1st equation such that it can be substituted into 2nd equation?

myininaya (myininaya):

\[(\frac{-1 +\sqrt{5}}{2})^3+2(\frac{-1+\sqrt{5}}{2})^2+2\] ---------- \[(\frac{-1 +\sqrt{5}}{2})^2=(\frac{-1 +\sqrt{5}}{2})(\frac{-1 + \sqrt{5}}{2})=\frac{1 -2 \sqrt{5}+5}{4}=\frac{6-2 \sqrt{5}}{4}\] ---------- \[(\frac{-1 + \sqrt{5}}{2})^3=(\frac{-1+\sqrt{5}}{2})^2(\frac{-1 + \sqrt{5}}{2})=(\frac{6 - 2 \sqrt{5}}{4})(\frac{-1 + \sqrt{5}}{2})\] \[=\frac{-6 +3 \sqrt{5}-2(5)}{8}=\frac{-16+3 \sqrt{5}}{8}\] ---------- \[(\frac{-1 +\sqrt{5}}{2})^3+2(\frac{-1+\sqrt{5}}{2})^2+2\] \[=\frac{-16 + 3 \sqrt{5}}{8}+2 \cdot \frac{6 - 2 \sqrt{5}}{4}+2\] \[=\frac{-16 + 3 \sqrt{5}}{8}+\frac{12- 4 \sqrt{5}}{4}+2\] \[=\frac{-16+3 \sqrt{5}+2(12-4 \sqrt{5})}{8}+2\] \[=\frac{8-5 \sqrt{5}}{8}+2=\frac{8-5 \sqrt{5}+8(2)}{8}=\frac{24-5 \sqrt{5}}{8}\] So I get not 3

myininaya (myininaya):

but that is for one of the values

myininaya (myininaya):

i leave the other one for you

OpenStudy (anonymous):

Sorry, I don't really get your reply...

myininaya (myininaya):

Well I found both of the values of x that solve x^2+x-1=0 Then I evaluated the expression x^3+2x^2+2 for one of the values I found (you can see what x I used above)